OpenStudy (anonymous):

Solve for h. h/5 - 16 = 8

pooja195 (pooja195):

add 16 to both sides

pooja195 (pooja195):

8+16=?

OpenStudy (anonymous):

Im writing it down toosoo

OpenStudy (anonymous):

24

OpenStudy (anonymous):

24

OpenStudy (anonymous):

h/5=24

pooja195 (pooja195):

Good not we need to multiply but the reciprocal so h/5 becomes 5/h or just 5 24 x 5 =?

pooja195 (pooja195):

*now

OpenStudy (anonymous):

120

OpenStudy (anonymous):

120

pooja195 (pooja195):

good so h=120